OpenStudy (anonymous):

solve. 3/2+3/4a=1/4a-1/2

OpenStudy (gw2011):

To make the problem easier multiply the entire equation by 4 to eliminate the denominators and you get: 6+3a=a-2 3a-a=-6-2 2a=-8 a=-4